Brianna's family spent $134 on 2 adults tickets and 3 youth tickets at an amusement park. Max's family spent $146 on 3 adult ticktes and 2 youth tickets. What is the price of a youth ticket?
Answer by richard1234(7193)

Let x be the price of an adult ticket and y be the price of a youth ticket. We have
2x + 3y = 134
3x + 2y = 146
Adding the two equations and dividing by 5, we obtain
x + y = 56
Since 2x + 3y = 2(x + y) + y = 2(56) + y = 112 + y we can substitute into the first equation to get
112 + y = 134 --> y = 22
Therefore the price of a youth ticket is $22.